OpenStudy (anonymous):

Use a addition or subtraction formula to simply for sin(3theta)*cos(theta)-cos(3theta)*sin(theta)=1/2, then find all solutions within interval [0, 2pi) (PS the simplified part is sin(2theta)=1/2)

OpenStudy (anonymous):

3theta and theta and it is cos(3theta-theta) which ends up simplifying to \[\sin(2\theta)=1/2\]

OpenStudy (aum):

\[ \sin(A-B) = \sin(A)\cos(B) - \cos(A)\sin(B) \]

OpenStudy (anonymous):

I need to find the correct solutions within the interval [0,2pi)

OpenStudy (aum):

\[ \sin(2\theta) = \frac 12 \\ 2\theta = \frac{\pi}{6}, \frac{5\pi}{6}, .... \]

OpenStudy (aum):

Add \(2\pi\) to the above to get two more angles. Then divide them all by 2 to get theta in the interval [0, 2pi]

OpenStudy (anonymous):

so the correct answers would be ... pi/12 and 5pi/12 within [0,2pi)

OpenStudy (aum):

\[\sin(2\theta) = \frac 12 \\ 2\theta = \frac{\pi}{6}, \frac{5\pi}{6}, \frac{\pi}{6} + 2\pi, \frac{5\pi}{6} + 2\pi \\ \theta = \frac{\pi}{12}, \frac{5\pi}{12}, \frac{\pi}{12} + \pi, \frac{5\pi}{12} + \pi \\ \theta = \frac{\pi}{12}, \frac{5\pi}{12}, \frac{13\pi}{12}, \frac{17\pi}{12} \]

OpenStudy (anonymous):

Thanks

OpenStudy (aum):

You are welcome.